OpenStudy (anonymous):

Help please! y=4mx+8b

OpenStudy (tyteen4a03):

What are you trying to do with this equation?

OpenStudy (anonymous):

Solve for x

OpenStudy (tyteen4a03):

Is this an linear equation? If so, is the slope 4 and the y-intercept 8?

OpenStudy (anonymous):

Its a literal equation. The answer must be one variable.

hartnn (hartnn):

subtract 8b from both sides of y=4mx+8b what u get ?

OpenStudy (anonymous):

-8b+y=4mx

hartnn (hartnn):

yes, now, divide both sides by 4m.

OpenStudy (anonymous):

ok but what do i do with the variables?

OpenStudy (tyteen4a03):

(Off-topic: If the slope is 4 and the y intercept 8, then the equation is y=4x+8, not y=4mx+8b. In this case, m is 4 and b is 8)

OpenStudy (anonymous):

The problem says 4mx.

hartnn (hartnn):

you can divide variable also...like if i tell you to divide ax+b=c by 3h, it would be (ax+b)/3h = c/3h so, what do you get after dividing 4m on both sides of -8b+y=4mx ??

OpenStudy (anonymous):

-8b+y/4m but how do i reduce that?

hartnn (hartnn):

\(x= \dfrac{-8b+y}{4m} = \dfrac{-8b}{4m}+\dfrac{y}{4m} \\ = \dfrac{-2b}{m}+\dfrac{y}{4m}\)

OpenStudy (anonymous):

Oh ok. But what do i do from there?

hartnn (hartnn):

i think you can leave your answer as (-8b+y)/4m and -2b/m + y/4m is most simplified...you cannot do anything further...

OpenStudy (anonymous):

Its not right.

hartnn (hartnn):

are you sure the question is y=4mx+8b ? if yes, x= (-8b+y)/4m must work... did you also try x= -2b/m + y/4m ?

OpenStudy (anonymous):

Yes. Im just gonna skip it because I cant figure it out.

OpenStudy (anonymous):

You were right. I just had to put it in a certain way. Thank you.

hartnn (hartnn):

welcome ^_^ glad i could help :)